# Measuring the Substructure Mass Power Spectrum of 23 SLACS Strong Galaxy-Galaxy Lenses with Convolutional Neural Networks

### v1

@@ Abstract
pubyear: 2023pagerange: Measuring the Substructure Mass Power Spectrum of 23 SLACS Strong Galaxy-Galaxy Lenses with Convolutional Neural NetworksB
@@ 1 Introduction
A common assumption in dark matter studies with lensing is treating its effect as a perturbation on a smooth lens potential. Although massive, isolated subhaloes can be modelled individually and have a localized effect, the much more common smaller mass subhaloes permeate the field of view and have a collective effect on the lensed features of the source. This has led many authors to treat such perturbations as a Gaussian Random Field (GRF, e.g. Chatterjee & Koopmans, 2017; Vernardos & Koopmans, 2022; Galan et al., 2022; Biggio et al., 2022; Bayer et al., 2023a, b). Alternatively one could use a parametric approach where dark matter substructure is modelled as an ensemble of individual subhaloes with Navarro–Frenk–White (NFW, Navarro et al., 1997) profiles (e.g. Wagner-Carena et al., 2022; Montel et al., 2022). In this work we adopt the statistical approach of treating the lens perturbation as a GRF parameterized by its power-law statistics.
Chatterjee & Koopmans (2017) first proposed a method for measuring the GRF power-law in simulated data by analysing the surface brightness anomalies in the lens, which are the residuals between the smooth lens model reconstruction and the data. They then predict the GRF power-law parameters through an MCMC analysis of the power spectrum. Bayer et al. (2023a) further developed the method of measuring the power spectrum of surface brightness anomalies and applied it to a sample of ten SLACS lenses, carefully modelling the noise properties of each lens to predict the power spectrum. As a pilot application, Bayer et al. (2023b) were able to obtain upper bound constraints on the GRF power spectrum for the single SLACS lens SDSS J0252+0039, selected due to its simple geometry and high signal-to-noise ratio. Recently Vernardos & Koopmans (2022); Galan et al. (2022); Biggio et al. (2022) also showed how a GRF perturbation to the lens potential could be recovered in mock data using three different techniques, but they have not yet been applied to real observations.
@@
We first present our selected sample of SLACS lenses in Sect. 2. We then describe the process of generating simulated strongly lensed images in Sect. 3, followed by the training of an uncertainty-aware CNN to predict the substructure mass power spectrum parameters in Sect. 4. We apply the trained CNNs to our SLACS lens sample and compare the results in Sect 5. We offer our concluding remarks and discussion in Sect. 6.
@@ 3 Lens simulation
In the absence of a large set of real observations, as is the case in strong lensing, a synthetic training set should be designed to be broad enough to include a large range of possible lenses but also include images that are very close to the observations. Vernardos et al. (2020) showed that the most important factor in this is the source brightness profile. We address this by using the reconstructed sources obtained by Shajib et al. (2021) to create training sets that are representative of our selected SLACS lens sample. In this section we present how we generate our training sets using these sources, as well as our assumptions on the lens mass, lens light, and instrumental effects.
@@ 3.1 Smooth lens potential
Table 1: Lens mass, lens light, and noise parameters and ranges for generating the mock data set. Each parameter is drawn uniformly within the given range. The lens light is modelled as an elliptical Sérsic profile with mean location, orientation, and axis ratio correlated to the lens mass model parameters $\{x_{0},y_{0},q,\theta_{\mathrm{pa}}\}$ but deviated by a Gaussian random variable with a standard deviation of $2.5$ per cent the maximum range (see Sect. 3.4). The effective radius $r_{\mathrm{eff}}$ of the Sérsic profile is proportional to the chosen Einstein radius $b$. The pixelated sources are described in Sect. 3.3. We also randomly select a point spread function to convolve with our simulated image from the 23 available (see Sect. 3.5).
@@ 3.3 Source brightness profile
We use the reconstructed sources of our sample of 23 SLACS lenses from Shajib et al. (2021) as source brightness profiles for our mock lenses. Details about their lens modelling technique can be found in section 3 of their paper, and the nominal source brightness profiles are shown in their figs. 2, 3, and 4. We focus on how they obtained their best-fitting sources as well as the full posterior distribution of the source parameters. The source galaxy brightness profiles are modelled using a basis of shapelets (Refregier, 2003) superimposed on an elliptical Sérsic profile (the elliptical Sérsic profile is described in Sect. 3.4). Shapelets are a complete set of orthogonal basis functions, and a brightness profile can be fit with a finite subset of the basis functions by optimizing the shapelet coefficients. This parameterization reduces the number of free source parameters by at least an order of magnitude (from thousands to less than a hundred) compared to using a pixelated grid (Birrer et al., 2015; Tagore & Jackson, 2016). Shapelet based source reconstructions have previously been used to model quadruply lensed quasars (Shajib et al., 2019) and time-delay lenses (Birrer et al., 2019; Shajib et al., 2020). Shajib et al. (2021) used MCMC sampling to obtain the posterior distributions of their model parameters, i.e. the lens mass parameters and the Sérsic and shapelet coefficients of the source.
@@
The correlated noise used for data augmentation is generating by convolving Gaussian white noise with a Gaussian kernel with standard deviation representing the length scale of the correlation, described in more detail in Sect. 3.5. An example of the data augmentation procedure is given in Fig. 3 for lens J1621+3931. After one augmentation the image remains very similar to the original best-fitting source (besides the rotation), however, with each subsequent augmentation it diverges further away.
@@
Our first method has the benefit of teaching the CNN to be invariant under small changes in the source morphology. In particular, the repeated multiplication of the nominal sources by correlated noise can change the power spectrum of the source images. We want to ensure the CNN is invariant to changes in the power spectrum of the sources to avoid confusing them with the power spectrum of the lensing perturbation. The downside is that we also include sources in our training set that can be far from the best-fitting. Our second method benefits from taking sources directly from the MCMC chain, which should be perturbed in a physically justifiable way and thus be more representative of the SLACS data. One potential caveat is that using only sources from the MCMC chain may systematically bias the training set, since the MCMC will not converge exactly to the true source brightness profile of the galaxy. We also create a third training set (the combined data set) by randomly selecting half of the prior two training sets. We later show in Sect. 5 that the resulting predictions on the SLACS lenses are consistent across CNNs trained with each data set, so the two methods perform similarly.
@@ 3.5 Point spread function and noise
We introduce correlated noise into our simulated images using a simpler and more general approach. We first create a grid of uncorrelated Gaussian white noise at the same resolution as the data and then convolve it with a Gaussian kernel to induce a correlation. The standard deviation of the Gaussian kernel, denoted $L_{\mathrm{cor}}$, determines the noise correlation length. Figure 5 shows examples of correlated noise for four different values of $L_{\mathrm{cor}}$. As $L_{\mathrm{cor}}$ increases, the noise gets more and more correlated between neighbouring pixels and appears clumpier. When $L_{\mathrm{cor}}=0^{\prime\prime}$, we recover the white noise case where each pixel is completely uncorrelated. We sample the correlation length uniformly in $L_{\mathrm{cor}}\in[0^{\prime\prime},0.05^{\prime\prime}]$ to induce Gaussian correlations with standard deviation up to one pixel length. Incorporating correlated noise into our simulation is particularly important, as the noise power spectrum could interfere with measuring the power spectrum of the GRF perturbation. It is essential for our CNNs to learn to distinguish between these two effects to obtain unbiased results.
The Gaussian noise level $\sigma_{\mathrm{noise}}$ (i.e. the root mean squared deviation of the noise) for each of our lenses was estimated by Shajib et al. (2021). Most of the SLACS lenses have very low noise levels, with J073+3216 being the noisiest with $\sigma_{\mathrm{noise}}=0.0048$ when the peak brightness is normalized to one. Most images have $\sigma_{\mathrm{noise}}<0.002$. We sample the logarithm of the noise level uniformly in $\log_{10}(\sigma_{\mathrm{noise}})\in[-3.25,-2.25]$ to encompass the SLACS data. We then create a realization of correlated noise for a given $\sigma_{\mathrm{noise}}$ and $L_{\mathrm{cor}}$, as described above, and add it to the lensed image. The images are normalized again by dividing by the peak brightness. This method of generating correlated noise is also used in our data augmentation scheme to generate source brightness profiles for data set 1, as described in Sect. 3.3.
@@ 3.6 Mock data set
The parameters required to generate a mock lens are listed in Table  1. Once a parameter vector is selected by uniformly drawing from the listed range, we use the Mock Lenses in Time software package [^3] https://github.com/gvernard/molet (MOLET, Vernardos, 2021) to simulate the images. Each data set consists of 250,000 lensed images with 230,000 used as the training set and 20,000 used as a validation set. The third combined data set uses a random half of data set 1 and 2. A random sample of 30 mock lensed images from the combined data set is shown in Appendix A.
@@ 4 Convolutional neural network
Our goal is to build and train a CNN to predict the GRF power spectrum parameters $\log_{10}(\sigma_{\delta\psi}^{2})$ and $\beta$. Uncertainty estimation plays a major role in our approach. In standard regression problems, a neural network is trained using supervised learning to produce point estimates. Only an overall systematic uncertainty can be estimated based on the network’s performance on the validation set using a metric such as the root mean squared error.
@@ 4.1.1 Training labels and aleatoric uncertainty
where we set $p_{0}=0.4$, $r=2.0$, and $n=0.9$. Fig. 6 shows how $p$ varies across the parameter space of $\log_{10}(\sigma_{\delta\psi}^{2})$ and $\beta$, matching the behaviour of Vernardos et al. (2020, fig. 3). Equation (6) is chosen to behave similarly to Vernardos et al. (2020, equation 4, 5) but only depend on the GRF power-law parameters. In Vernardos et al. (2020) $n=0.6$ was arbitrarily used, while here we use a larger value to avoid underestimating the uncertainty (further discussed in Sect. 4.3). By design, the ground truth value could be anywhere in the uniform probability distributions of the training labels but tends to closer to the centre. Figure 7 shows an example target and predicted distribution of our fully trained CNN for one mock image in our validation set. After every epoch the training labels are regenerated, drawing new values from the binomial distribution each time.
@@
Figure 7: Target distributions (dashed orange lines) and predictions of the CNN (blue solid lines) for $\log_{10}(\sigma_{\delta\psi}^{2})$ (top) and $\beta$ (bottom) for one example mock image of the validation set of the combined data set. The true value is indicated by the vertical dashed line but is never given directly to the CNN. Our predictions are obtained by drawing 200 samples from our BNN and averaging them. The first five samples are shown in Fig. 8.
@@ 4.1.2 Epistemic uncertainty
Figure 8: Example of multiple predictions made by our BNN for the same validation example as in Fig. 7. The first five samples are shown in different colours for $\log_{10}(\sigma_{\delta\psi}^{2})$ (top) and $\beta$ (bottom). The true value is indicated by the dashed vertical line. We could use the variation in the mean as a measure of the epistemic uncertainty, but in our case, we are only concerned with the combined aleatoric and epistemic uncertainties, i.e. the average probability distribution across many samples.
@@
We use a Bayesian Neural Network (BNN) to obtain uncertainty in the model parameters. Each time the network makes a prediction, a random value of each weight is drawn from a Gaussian probability distribution with trainable mean and variance rather than having fixed weights across the network (Graves, 2011). During inference, we repeatedly sample from the network to obtain an ensemble of different predictions. The epistemic uncertainty is included by averaging across the ensemble. Figure 8 shows five predictions of the network for the same validation example as in Fig. 7. The epistemic uncertainty can be represented as the variation in the mean prediction across the different samples. The aleatoric and epistemic uncertainties are combined by averaging across 200 different predictions from the BNN.
@@
Several methods are available for efficiently sampling from the weight distributions. We use the most common one, Flipout (Wen et al., 2018), an efficient technique for optimizing a BNN by generating pseudo-independent weight perturbations on mini-batches of data. The log-variance of the kernel posterior is initialized by $\mathcal{N}(-9,0.01)$ consistent with Hortúa et al. (2020). In variational inference, the negative evidence lower bound object (ELBO) is minimized with respect to the variational parameters (Graves, 2011; Gal & Ghahramani, 2016; Perreault Levasseur et al., 2017). The Kullback–Leibler (KL) divergence between the variational distribution and the prior is included in our loss to act as a regularization term to control the standard deviation of each weight in the network. We also use KL-annealing (Fu et al., 2019) linearly for the first 100 epochs to slowly introduce the KL-divergence term. The authors of Hortúa et al. (2020) discuss how $L_{2}$ regularization can be used on the standard deviation of each weight to calibrate the uncertainty in our predictions. We apply an $L_{2}$ regularization of $10^{-3}$ on the standard deviation. Lowering the regularization leads to higher uncertainties on each weight, while increasing the regularization makes the network behave more deterministic by pushing the variances closer to zero. This parameter can be adjusted in the same way that the dropout rate can be adjusted using Monte Carlo dropout (Gal & Ghahramani, 2016) to calibrate the uncertainty (Perreault Levasseur et al., 2017). We note that the BNN will have twice as many parameters as a corresponding deterministic network because we optimize both the mean and standard deviation of each weight.

### v2

@@ Abstract
pubyear: 2023pagerange: Measuring the Substructure Mass Power Spectrum of 23 SLACS Strong Galaxy-Galaxy Lenses with Convolutional Neural Networks22
@@ 1 Introduction
A common assumption in dark matter studies with lensing is treating its effect as a perturbation on a smooth lens potential. Although massive, isolated subhaloes can be modelled individually and have a localized effect, the much more common smaller mass subhaloes permeate the field of view and have a collective effect on the lensed features of the source. This has led many authors to treat such perturbations as a Gaussian Random Field (GRF; Chatterjee & Koopmans, 2017; Vernardos & Koopmans, 2022; Galan et al., 2022; Biggio et al., 2022; Bayer et al., 2023a, b). Alternatively one could use a parametric approach where dark matter substructure is modelled as an ensemble of individual subhaloes with Navarro–Frenk–White (NFW; Navarro et al., 1997) profiles (e.g., Wagner-Carena et al., 2022; Montel et al., 2022). In this work we adopt the statistical approach of treating the lens perturbation as a GRF parameterized by its power-law statistics.
Chatterjee & Koopmans (2017) first proposed a method for measuring the GRF power-law in simulated data by analysing the surface brightness anomalies in the lens, which are the residuals between the smooth lens model reconstruction and the data. They then predict the GRF power-law parameters through an MCMC analysis of the power spectrum. Bayer et al. (2023a) further developed the method of measuring the power spectrum of surface brightness anomalies and applied it to a sample of ten SLACS lenses, carefully modelling the noise properties of each lens to predict the power spectrum. As a pilot application, Bayer et al. (2023b) were able to obtain upper bound constraints on the GRF power spectrum for the single SLACS lens SDSS J0252+0039, selected due to its simple geometry and high signal-to-noise ratio. Recently, Vernardos & Koopmans (2022); Galan et al. (2022); Biggio et al. (2022) also showed how a GRF perturbation to the lens potential could be recovered in mock data using three different techniques, but they have not yet been applied to real observations.
@@
We first present our selected sample of SLACS lenses in Section 2. We then describe the process of generating simulated strongly lensed images in Section 3, followed by the training of an uncertainty-aware CNN to predict the substructure mass power spectrum parameters in Section 4. We apply the trained CNNs to our SLACS lens sample and compare the results in Sect 5. We offer our concluding remarks and discussion in Section 6.
@@ 3 Lens simulation
In the absence of a large set of real observations, as is the case in strong lensing, a synthetic training set should be designed to be broad enough to include a large range of possible lenses but also include images that are very close to the observations. Vernardos et al. (2020) showed that the most important factor in this is the source brightness profile. We address this by using the reconstructed sources obtained by Shajib et al. (2021) to create training sets that are representative of our SLACS lens sample. In this section we present how we generate our training sets using these sources, as well as our assumptions on the lens mass, lens light, and instrumental effects.
@@ 3.1 Smooth lens potential
Table 1: Lens mass, lens light, and noise parameters and ranges for generating the mock data set. Each parameter is drawn uniformly within the given range. The lens light is modelled as an elliptical Sérsic profile with mean location, orientation, and axis ratio correlated to the lens mass model parameters $\{x_{0},y_{0},q,\theta_{\mathrm{pa}}\}$ but deviated by a Gaussian random variable with a standard deviation of $2.5$ per cent the maximum range (see Section 3.4). The effective radius $r_{\mathrm{eff}}$ of the Sérsic profile is proportional to the chosen Einstein radius $b$. The pixelated sources are described in Section 3.3. We also randomly select a point spread function to convolve with our simulated image from the 23 available (see Section 3.5).
@@ 3.3 Source brightness profile
We use the reconstructed sources of our sample of 23 SLACS lenses from Shajib et al. (2021) as source brightness profiles for our mock lenses. Details about their lens modelling technique can be found in section 3 of their paper, and the nominal source brightness profiles are shown in their figs. 2, 3, and 4. We focus on how they obtained their best-fitting sources as well as the full posterior distribution of the source parameters. The source galaxy brightness profiles are modelled using a basis of shapelets (Refregier, 2003) superimposed on an elliptical Sérsic profile (the elliptical Sérsic profile is described in Section 3.4). Shapelets are a complete set of orthogonal basis functions, and a brightness profile can be fit with a finite subset of the basis functions by optimizing the shapelet coefficients. This parameterization reduces the number of free source parameters by at least an order of magnitude (from thousands to less than a hundred) compared to using a pixelated grid (Birrer et al., 2015; Tagore & Jackson, 2016). Shapelet based source reconstructions have previously been used to model quadruply lensed quasars (Shajib et al., 2019) and time-delay lenses (Birrer et al., 2019; Shajib et al., 2020). Shajib et al. (2021) used MCMC sampling to obtain the posterior distributions of their model parameters, i.e. the lens mass parameters and the Sérsic and shapelet coefficients of the source.
@@
The correlated noise used for data augmentation is generating by convolving Gaussian white noise with a Gaussian kernel with standard deviation representing the length scale of the correlation, described in more detail in Section 3.5. An example of the data augmentation procedure is given in Fig. 3 for lens J1621+3931. After one augmentation the image remains very similar to the original best-fitting source (besides the rotation), however, with each subsequent augmentation it diverges further away.
@@
Our first method has the benefit of teaching the CNN to be invariant under small changes in the source morphology. In particular, the repeated multiplication of the nominal sources by correlated noise can change the power spectrum of the source images. We want to ensure the CNN is invariant to changes in the power spectrum of the sources to avoid confusing them with the power spectrum of the lensing perturbation. The downside is that we also include sources in our training set that can be far from the best-fitting. Our second method benefits from taking sources directly from the MCMC chain, which should be perturbed in a physically justifiable way and thus be more representative of the SLACS data. One potential caveat is that using only sources from the MCMC chain may systematically bias the training set, since the MCMC will not converge exactly to the true source brightness profile of the galaxy. We also create a third training set (the combined data set) by randomly selecting half of the prior two training sets. We later show in Section 5 that the resulting predictions on the SLACS lenses are consistent across CNNs trained with each data set, so the two methods perform similarly.
@@ 3.5 Point spread function and noise
We introduce correlated noise into our simulated images using a simpler and more general approach. We first create a grid of uncorrelated Gaussian white noise at the same resolution as the data and then convolve it with a Gaussian kernel to induce a correlation. The standard deviation of the Gaussian kernel, denoted $L_{\mathrm{cor}}$, determines the noise correlation length. Figure 5 shows examples of correlated noise for four different values of $L_{\mathrm{cor}}$. As $L_{\mathrm{cor}}$ increases, the noise gets more and more correlated between neighbouring pixels and appears clumpier. When $L_{\mathrm{cor}}=0^{\prime\prime}$, we recover the white noise case where each pixel is completely uncorrelated. We sample the correlation length uniformly in $L_{\mathrm{cor}}\in[0^{\prime\prime},0.05^{\prime\prime}]$ to induce Gaussian correlations with a standard deviation up to one pixel length. Incorporating correlated noise into our simulation is particularly important, as the noise power spectrum could interfere with measuring the power spectrum of the GRF perturbation. It is essential for our CNNs to learn to distinguish between these two effects to obtain unbiased results.
The Gaussian noise level $\sigma_{\mathrm{noise}}$ (i.e. the root mean squared deviation of the noise) for each of our lenses was estimated by Shajib et al. (2021). Most of the SLACS lenses have very low noise levels, with J073+3216 being the noisiest with $\sigma_{\mathrm{noise}}=0.0048$ when the peak brightness is normalized to one. Most images have $\sigma_{\mathrm{noise}}<0.002$. We sample the logarithm of the noise level uniformly in $\log_{10}(\sigma_{\mathrm{noise}})\in[-3.25,-2.25]$ to encompass the SLACS data. We then create a realization of correlated noise for a given $\sigma_{\mathrm{noise}}$ and $L_{\mathrm{cor}}$, as described above, and add it to the lensed image. The images are normalized again by dividing by the peak brightness. This method of generating correlated noise is also used in our data augmentation scheme to generate source brightness profiles for data set 1, as described in Section 3.3.
@@ 3.6 Mock data set
The parameters required to generate a mock lens are listed in Table  1. Once a parameter vector is selected by uniformly drawing from the listed range, we use the Mock Lenses in Time software package [^3] https://github.com/gvernard/molet (MOLET; Vernardos, 2021) to simulate the images. Each data set consists of 250,000 lensed images with 230,000 used as the training set and 20,000 used as a validation set. We also include a final test set for each method with 25,000 images. The third combined data set uses a random half of data set 1 and 2. A random sample of 30 mock lensed images from the combined data set is shown in Appendix A.
@@ 4 Convolutional neural network
Our goal is to build and train a CNN to predict the GRF power spectrum parameters $\log_{10}(\sigma_{\delta\psi}^{2})$ and $\beta$. Uncertainty estimation plays a major role in our approach. In standard regression problems, a neural network is trained using supervised learning to produce point estimates. Only an overall systematic uncertainty can be estimated based on the network’s performance on the test set using a metric such as the root mean squared error.
@@ 4.1.1 Training labels and aleatoric uncertainty
where we set $p_{0}=0.4$, $r=2.0$, and $n=0.9$. Fig. 6 shows how $p$ varies across the parameter space of $\log_{10}(\sigma_{\delta\psi}^{2})$ and $\beta$, matching the behaviour of Vernardos et al. (2020, fig. 3). Equation (6) is chosen to behave similarly to Vernardos et al. (2020, equation 4, 5) but only depend on the GRF power-law parameters. In Vernardos et al. (2020) $n=0.6$ was arbitrarily used, while here we use a larger value to avoid underestimating the uncertainty (further discussed in Section 4.3). By design, the ground truth value could be anywhere in the uniform probability distributions of the training labels but tends to closer to the centre. Figure 7 shows an example target and predicted distribution of our fully trained CNN for one mock image in our test set. After every epoch the training labels are regenerated, drawing new values from the binomial distribution each time.
@@
Figure 7: Target distributions (dashed orange lines) and predictions of the CNN (blue solid lines) for $\log_{10}(\sigma_{\delta\psi}^{2})$ (top) and $\beta$ (bottom) for one example mock image of the test set of the combined data set. The true value is indicated by the vertical dashed line but is never given directly to the CNN. Our predictions are obtained by drawing 200 samples from our BNN and averaging them. The first five samples are shown in Fig. 8.
@@ 4.1.2 Epistemic uncertainty
Figure 8: Example of multiple predictions made by our BNN for the same test set example as in Fig. 7. The first five samples are shown in different colours for $\log_{10}(\sigma_{\delta\psi}^{2})$ (top) and $\beta$ (bottom). The true value is indicated by the dashed vertical line. We include the epistemic uncertainty by averaging across the different samples, increase the overall uncertainty by widening the probability distributions.
@@
We use a Bayesian neural network (BNN) to obtain uncertainty in the model parameters. Each time the network makes a prediction, a random value of each weight is drawn from a Gaussian probability distribution with trainable mean and variance rather than having fixed weights across the network (Graves, 2011). During inference, we repeatedly sample from the network to obtain an ensemble of different predictions. The epistemic uncertainty is included by averaging across the ensemble of 200 predictions from the BNN. Figure 8 shows five predictions of the network for the same test set example as in Fig. 7. The epistemic uncertainty can be represented as the increase in the uncertainty due to averaging across the different samples.
@@
Several methods are available for efficiently sampling from the weight distributions of a BNN. We use the most common one, Flipout (Wen et al., 2018), an efficient technique for optimizing a BNN by generating pseudo-independent weight perturbations on mini-batches of data. Unlike traditional methods where weight samples are shared across all examples in a mini-batch, Flipout applies a random sign matrix to the weights for each data point. This approach efficiently decorrelates the gradients between different examples in the mini-batch, reducing the gradient variance compared with shared perturbations and thereby stabilizing the training process. Hortúa et al. (2020) showed that Flipout outperforms other methods of sampling from the weight distributions of a BNN during training, including Dropout, DropConnect, and the reparameterization trick. The log-variance of the kernel posterior is initialized by $\mathcal{N}(-9,0.01)$ consistent with Hortúa et al. (2020). In variational inference, the negative evidence lower bound object (ELBO) is minimized with respect to the variational parameters (Graves, 2011; Gal & Ghahramani, 2016; Perreault Levasseur et al., 2017). The Kullback–Leibler (KL) divergence between the variational distribution and the prior is included in our loss to act as a regularization term to control the standard deviation of each weight in the network. We also use KL-annealing (Fu et al., 2019) linearly for the first 100 epochs to slowly introduce the KL-divergence term. The authors of Hortúa et al. (2020) discuss how $L_{2}$ regularization can be used on the standard deviation of each weight to calibrate the uncertainty in our predictions. We apply an $L_{2}$ regularization of $10^{-3}$ on the standard deviation. Lowering the regularization leads to higher uncertainties on each weight, while increasing the regularization makes the network behave more deterministic by pushing the variances closer to zero. This parameter can be adjusted in the same way that the dropout rate can be adjusted using Monte Carlo dropout (Gal & Ghahramani, 2016) to calibrate the uncertainty (Perreault Levasseur et al., 2017). We note that the BNN will have twice as many parameters as a corresponding deterministic network because we optimize both the mean and standard deviation of each weight.